# HumanoidBench: Simulated Humanoid Benchmark for Whole-Body Locomotion and Manipulation

### v1

@@ V-E Common Failures
Common Failure on highbar. In the highbar task, the Unitree H1 robot conservatively learns to maintain contact with the bar to avoid episode termination, but experiences difficulties in performing the whole-body rotation trajectory. This is indicative of short horizon planning and, despite the availability of dense rewards, is a recurrent challenge in many of the long-horizon benchmark tasks.
@@ VI Conclusion
Here we have focused on reinforcement learning algorithms because humanoids are a setting where collecting physical demonstrations may be challenging. However, we believe that other means could be employed to bootstrap learning (e.g., learning from human videos).
@@ Appendix A Additional Components
In addition to the Unitree H1 robot, we also include model files for the Agility Robotics Digit humanoid. We add a custom head, together with egocentric vision and whole-body tactile sensing, similarly to the H1. We provide a torque controlled version of this robot.
We also provide two additional end effectors for the Unitree H1 robot, that is, a Robotiq 2F-45 parallel-jaw gripper, and a dexterous hand available in the Unitree model collection.
@@ B-A Observation Space
The observation space for the robot state (joint positions and velocities) comprises $151$ dimensions, with $49$ dimensions representing the humanoid robot body and $51$ dimensions representing each hand. The observation space can also include environment states for tasks interacting with objects, as described in detail in Section B-E.
@@ B-C Simulation Performance
HumanoidBench is based on MuJoCo [52], which provides fast and accurate physics simulation. We provide a diverse set of configurations, such as the full humanoid model with two hands, the humanoid model without hands, and the humanoid model with collision meshes only on its feet. We use the fastest model (i.e. collision meshes only on feet) to train the low-level reaching policies, described in Section C-B.
@@
TABLE II: HumanoidBench Simulation Performance.
@@ B-E Task Specification
Objective. Open four different types of cabinet doors (e.g. hinge door, sliding door, drawer) and perform different manipulations (see subtasks below) for objects inside the cabinet.
@@ C-C Benchmarking Results
TABLE III: Average returns for HumanoidBench. Each number represents average return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).
@@
TABLE IV: Maximum returns for HumanoidBench. Each number represents maximum return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).

### v2

@@ V-E Common Failures
Common Failure on highbar. In the highbar task, the Unitree H1 robot conservatively learns to maintain contact with the bar to avoid episode termination, but experiences difficulties in performing the whole-body rotation trajectory. This is indicative of short horizon planning and a recurrent challenge in many of the long-horizon benchmark tasks, despite the availability of dense rewards.
@@ VI Conclusion
Here we have focused on reinforcement learning algorithms because collecting physical demonstrations with humanoid robots is particularly challenging. However, we believe that other means could be employed to bootstrap learning (e.g., learning from human videos).
@@
Acknowledgments
This work was supported in part by the SNSF Postdoc Mobility Fellowship 211086, ONR MURI N00014-22-1-2773, BAIR Industrial Consortium, Komatsu, InnoHK Centre for Logistics Robotics, an ONR DURIP grant, the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University) and RS-2024-00333634). We also thank Google TPU Research Cloud (TRC) for granting us access to TPUs for research.
@@ Appendix A Additional Components
(a) Digit w/ Shadow Hands
(b) Unitree G1
(c) H1 w/ Robotiq grippers
(d) H1 w/ Unitree hands
Figure 9: Additional configurations available in HumanoidBench.
In addition to the Unitree H1 robot, we also include model files for the Agility Robotics Digit humanoid (see Figure 9). We add a custom head, together with egocentric vision and whole-body tactile sensing, similarly to the H1. We provide a torque controlled version of this robot. We also provide a torque controlled version of the more recent Unitree G1 humanoid.
For the broader use of HumanoidBench, we also provide two simple end-effector options for the Unitree H1 robot: a $2$-DoF Robotiq 2F-45 parallel-jaw gripper (with a rotational wrist joint), and a $13$-DoF dexterous Unitree hand available in the Unitree model collection.Footnote 5
Specifications for all these variants are detailed in Table III.
H1 w/o hands H1 w/ ShadowHand H1 w/ Robotiq gripper H1 w/ Unitree hand Digit w/ ShadowHand Unitree G1
Observation space $51$ $151$ $55$ $103$ $221$ $87$
Action space $19$ $61$ $23$ $45$ $65$ $37$
DoF (body) $25$ $25$ $25$ $25$ $57$ $29$
DoF ($2$ end-effectors) 00 $50$ $4$ $26$ $50$ $14$
TABLE III: All supported robot specifications. Note that the observation space in this table does not take into account any observations of the surrounding environment and solely comprises generalized positions and velocities. We use quaternions for the robot floating base orientation, as well as for ball joints, which add additional position coordinates compared to the velocity components (which match the DoFs).
@@ B-A Observation Space
The observation space for the benchmarked H1 robot state (joint positions and velocities) comprises $151$ dimensions, with $51$ dimensions representing the humanoid robot body and $50$ dimensions representing each hand. Table III summarizes observation spaces for other robotic configurations supported in HumanoidBench. The observation space can also include environment states for tasks interacting with objects, as described in detail in Section B-E.
@@ B-C Simulation Performance
HumanoidBench is based on MuJoCo [60], which provides fast and accurate physics simulation. We provide a diverse set of configurations, such as the full humanoid model with two hands, the humanoid model without hands, and the humanoid model with collision meshes only on its feet. We use the fastest model (i.e., collision meshes only on feet) to train the low-level reaching policies, described in Section C-B.
@@
TABLE IV: HumanoidBench Simulation Performance.
@@ B-E Task Specification
Objective. Open four different types of cabinet doors (e.g., hinge door, sliding door, drawer) and perform different manipulations (see subtasks below) for objects inside the cabinet.
@@ C-C Benchmarking Results
TABLE V: Average returns for HumanoidBench. Each number represents average return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).
@@
TABLE VI: Maximum returns for HumanoidBench. Each number represents maximum return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).